# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v2

@@ 1. Introduction
In recent years, technological advancements in Generative Artificial Intelligence (Goodfellow et al., 2014; Li et al., 2022; Touvron et al., 2023; Alayrac et al., 2022) has been driven by a class of transformer-based (Vaswani et al., 2017) models called foundational models (FMs) (Awais et al., 2023; Bansal et al., 2024). FMs, such as large language models (LLMs, e.g. Meta’s LLaMA3 (Meta, 2024)) and vision language models (VLMs, e.g. Microsoft’s LLaVA (Liu et al., 2024a))  [^1] In this paper the term FMs is used interchangeably with LLMs and VLMs. are billion-scale parameter models, trained on a large corpora of data. FM scaling laws (Zhang et al., 2024; Bi et al., 2024; Bambhaniya et al., 2024) correlate better learning with model size (Yang et al., 2024). Thus, inference with FMs has high demand for memory, data movement, energy, and computation.
Recent research has focused on various model compression techniques such as pruning (Frantar and Alistarh, 2023; Ashkboos et al., 2024a) and quantization (Lin et al., 2024a; Shao et al., 2023; Guo et al., 2023; Kang et al., 2024) to reduce memory footprint and computational costs, to make FM inference service more accessible to resource-constrained devices.
Model pruning (Frantar and Alistarh, 2023) reduces memory footprint by removing ineffectual model parameters, such as individual weights (unstructured) or blocks of weights (structured), and storing sparse tensors in a compressed format (Jeong et al., 2023). However, pruning of FMs may be infeasible due to, significant accuracy drops even at low pruning ratios (Kuzmin et al., 2024; Yin et al., [n. d.]), with potential demand for compute and memory-intensive fine-tuning to regain accuracy. Model quantization, on the other hand, reduces the size of a target model by representing weights and/or activations at low precision (Ramachandran et al., 2024b; Kang et al., 2024; Ramachandran et al., 2024a). Quantization not only reduces the memory footprint but also enables efficient low-bit-width arithmetic operations (You et al., 2024; van Baalen et al., 2023). Recent works on quantization (Dettmers et al., 2022; Guo et al., 2023) have identified that quantizing LLMs is considerably more challenging than quantizing traditional DNNs (Li et al., 2021; Tambe et al., 2020) due to the emergence of large magnitude features known as outliers (Zadeh et al., 2020). These outliers significantly impact model accuracy and require specialized handling (Guo et al., 2023) compared to inliers (small magnitude features).
Table 1. Comparison of MicroScopiQ with prior outlier-aware co-design techniques in groups A (Zadeh et al., 2020; Park et al., 2018) and B (Guo et al., 2023).
@@ Methods
Techniques in group A, such as OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2023) (algorithm) and GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) (architecture co-design) exhibit low accuracy degradation. This is because they typically store outliers at high precision separated from lower precision inliers. However, these techniques result in, (a) low compression factor with high effective bit-width (EBW [^2] The average number of bits used to represent each quantized parameter of a model.) and (b) inefficient hardware and unaligned memory access.
@@
Based on the limitations of existing solutions discussed above, we identify that assignment of higher bit-width for outliers is essential for good accuracy while for aligned memory and hardware efficiency a consistent bit-budget 333number of bits per weight/activation element and data type per tensor element is desired. Here, by consistency we mean that on average each scalar should be represented by a fixed bit-width of a particular data-type. However, these demands are generally conflicting to each each other.
@@
Towards achieving this, we present a novel co-design technique for the post-training weight quantization (PTQ) of FMs, namely MicroScopiQ. Our approach effectively leverages pruning with outlier aware quantization to achieve both memory alignment and improved accuracy. To effectively perform this, for a layer we prune the least important weights based on the Hessian information, and redistribute the additional bits of the outlier weights in the pruned locations. This ensures memory alignment while allowing outlier weights to have higher precision. Additionally, to maintain low error we use the recently proposed MicroScaling (MX) FP data format (Rouhani et al., 2023; Darvish Rouhani et al., 2023) to quantize outlier weights compared to MX-INT inlier quantization. To efficiently support outliers in a different format with different bits at different locations, we present an intelligent NoC architecture called, Redistribution and Coordination NoC (ReCoN). It offers minimal overhead and high throughput outlier processing and reorganization. We then present an accelerator that leverages ReCoN with a simple, homogeneous INT-PE array. Further, we extend the accelerator to be generic enough to support multiple bit-precision (2/4-bit operations). As summarized in Table 1, MicroScopiQ blends the advantages of group A and B techniques while mitigating their specific drawbacks.
@@ 2.1. Model Quantization
where, $\emph{max}_{\tau}^{b}$ is the maximum representable value of a data format (Ramachandran et al., 2024a). For $b$-bit INT quantization, $\emph{max}_{INT}^{b}=2^{b-1}-1$. After determining the quantization parameters, the quantized tensor can be calculated as (Ramachandran et al., 2024a),
@@
In model quantization, the quantization parameters can be shared at different granularities for different accuracy-overhead trade-offs. In increasing order of overheads, we have per-tensor quantization, wherein the scale factor is shared among all tensor elements. In per-channel quantization, the scale factor is shared per row/column of a tensor. Finally, in group quantization, the parameters are shared at a finer granularity between groups of $k$ (64, 128 etc.) elements in a row or column. These groups are formed by dividing channels into multiple non-overlapping contiguous blocks. In this paper, we adopt MX-FP and MX-INT quantization for inliers and outliers, respectively.
@@ 2.2. Microscaling Data Format
The MX data format proposed by prior works (Rouhani et al., 2023; Darvish Rouhani et al., 2023), is standardized by the Open Compute Project (Project, 2024) with support from Microsoft, Intel, NVIDIA and others. As shown in Figure 1, MX is a variant of block data representation (BDR) (Darvish Rouhani et al., 2023) that defines a format to represent a group of values collectively using shared scale factors. It leverages multi-level, power-of-two scaling, at fine- (level-1, $k_{1}$) and ultra-fine (level-2, $k_{2}$) granularity (Dai et al., 2021; Drumond et al., 2018). The MX data format is characterized by four components: i) scale factors (level-1, 2), ii) data type ($\tau$), iii) bit-width ($b$) and iv) group sizes ($k_{1},k_{2}$). In this paper we denote an MX-FP format as MX-FP-$b$${}_{k_{1},k_{2}}$. In this work, we adopt the version of the MX-FP data format proposed in (Rouhani et al., 2023) , employing multi-level scaling, which differs from the single-level scaling used in the standardized format (Project, 2024). The level-1 scale factor for MX-FP is computed following Equation 1. Conversely, for level-2 scale factor, we identify that the MX-FP leverages the sharing of exponent field of FP values (Wang et al., 2018) (referred as $\mu X$ in Figure 1). We show later in section 4.2 that by taking advantage of this insight i.e., the concept of shared $\mu X$, we are able to represent FP-outliers in INT format, thereby, enabling the design of simple, homogeneous INT-based PEs. For inliers, we employ MX-INT-$b$${}_{k_{1}}$ with a single level of scale factor following (Project, 2024). This is because, INT format does not possess an exponent field, thereby, a level-2 scale factor similar to MX-FP is not applicable. For simplified understanding, MX-INT-$b$${}_{k_{1}}$ inlier quantization can be viewed as analogous to INT group quantization utilizing an E8M0 scale factor. We refer the interested reader to (Darvish Rouhani et al., 2023) for an in-depth evaluation of the MX format and comparison against prior BDR formats.
Figure 2. ((a) Layer-wise distribution of outliers and adjacent outliers for different LLMs and VLMs as a percentage of total number of weights, (b) Weight-only quantization accuracy comparison between OliVe-W4 and MicroScopiQ-W2 for the above models on various benchmarks ($\text{Model}_{\text{benchmark}}$).
@@ 3.1. Limitations of existing techniques
In Table 1, we compare candidate proposals from group A: GOBO (Zadeh et al., 2020) and group B: OliVe (Guo et al., 2023) across various metrics. GOBO is able to achieve high accuracy by retaining outliers at full-precision. It stores outliers separately from low-precision inliers by using sparse representations with the associated outlier indices (see Figure 3(b)). By retaining outliers at full-precision, GOBO results in high EBW. Moreover, the compressed sparse storage and multiple precisions results in unaligned and random memory accesses (Guo et al., 2023), significantly impacting inference latency. Furthermore, GOBO’s outlier handling is hardware inefficient, requiring complex PE and control, with significant hardware overheads.
OliVe (Guo et al., 2023) proposes a scheme to ensure aligned memory accesses by quantizing inliers and outliers at the same precision (low EBW), but using different data formats. To enable differentiation between the inlier and outlier formats, it prunes the value adjacent to the outlier for use as an identifier (see Figure 3(c)). However, OliVe results in significant accuracy degradation, especially at low precision (see Figure 2(b)), due to: 1) sacrificing a number encoding from inliers for exclusive use as an identifier, reducing the number of representable values in the quantized range, and 2) the flawed assumption of outlier locality–particularly for modern FMs–that outliers are never adjacent to each other and only inliers are almost always adjacent to outliers (see section 3.2), leading to unintended outlier pruning. Furthermore, it still requires a fairly complex PE design incurring significant encoding/decoding overheads to convert the different formats into a unified processing format (exponent-integer pair) . In this paper, we show that despite quantizing outliers at higher-precision and in a different format, we ensure aligned memory access, simple PE design and minimal hardware overhead.
@@ 3.2. Adjacent Outliers Matter
Similar to prior works (Guo et al., 2023; Park et al., 2018), we leverage the $3\sigma$ rule (Pukelsheim, 1994) to categorize weights as outliers. We visually demonstrate the distribution of outliers and adjacent outliers 444pair of contiguous outliers as a percentage of the total number of weights in a layer across different FMs in Figure 2(a). As the orange box-plot shows, outliers depict a maximum percentage of $\sim 5.1\%$. Outliers are prevalent in FMs, and preserving their values is crucial for maintaining quantized model accuracy. Importantly, from the green box-plots, we observe that modern FMs on average possess $>0.5\%$ adjacent outliers per layer, with some FM layers showing peaks of $>2\%$. This is in stark contrast to the models evaluated by OliVe, such as BERT (Devlin et al., 2018) and OPT (Zhang et al., 2022) which have $<0.04\%$ adjacent outliers (two orders of magnitude lower than FMs like LLaMA3 and LlaVa). This indicates that while pruning values adjacent to outliers could have been ideal for models like BERT (Devlin et al., 2018), it is sub-optimal for most modern FMs as it removes crucial outlier values, leading to higher accuracy degradation. This is evident from Figure 2(b) where OliVe has significant accuracy degradation at 4-bit quantization due to its assumption of outlier locality. Unlike OliVe, MicroScopiQ does not naively prune adjacent values; instead it leverages the Hessian information (Frantar et al., 2022) to identify the least important values to prune ensuring outliers are preserved. This directly translates to high quantized model accuracy and MicroScopiQ at 2-bit weight quantization consistently outperforms OliVe across different FMs.
@@ 3.3. Outlier Precision and Data Format
The ability of group A techniques like GOBO (Zadeh et al., 2020) to achieve high quantized model accuracy even at extreme quantization levels of inliers ($<4$-bits) is due to retaining outliers at higher precision. This is particularly crucial at ultra-low bit width quantization because, if inliers and outliers are to be quantized to the same precision, there will be higher outlier quantization error due to the reduced representational range. We demonstrate this effect on the MicroScopiQ quantized FM accuracy in Table 6 wherein the quantized FM has poor performance when inliers and outliers are at 2-bits compared to outliers at 4-bits. Furthermore, evidence from recent work (Wu et al., 2023) demonstrates that FP-based formats for LLMs results in superior quantization performance compared to INTs. To validate this, we compare INT v/s MX-FP inlier and outlier quantization in Table 6. Evidently using MX-FP instead of INT for outliers results in better performance. The performance boost from employing MX-FP is due to higher dynamic range of FPs, which is particularly beneficial at extreme quantization levels. In this work, we quantize outliers at a higher precision (2$\times$) compared to inliers, using MX-FP for outliers and MX-INT for inliers.
@@ 4. MicroScopiQ Quantization Methodology
We present an overview of the MicroScopiQ quantization framework in Figure 3(a) and detail its steps in Algorithm 1. MicroScopiQ supports various group size granularities and any inlier and outlier (2×2$\times$2 × inliers) data precision. For simplicity, we explain with inlier precision of 2- and 4-bit and group sizes of 128 for inliers and 8 (level-1 and 2) for outliers. Experimental results in section 6 demonstrate the effects of different precisions and group sizes.
Figure 3. (a) Overview of the proposed MicroScopiQ quantization framework depicting methodology of inlier and outlier quantization and redistribution of outlier bits for a sample LLM weight matrix. Comparison against prior quantization frameworks (b) GOBO, and (c) OliVe
@@ 4.1. Preliminaries
We evaluate the second order derivative of Equation 3, namely the Hessian (Hassibi et al., 1993; LeCun et al., 1989) through Taylor series expansion (Hassibi et al., 1993). Note, the Hessian is the same for all rows depending only on input data and is given as, $\mathbb{H}=2\mathbb{X}\mathbb{X}^{T}$. Its inverse is, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$. We leverage the Hessian information to identify weights with small “saliencies” for pruning (L17 in Algo. algorithm 1). The number of inliers to prune is determined by the number of outliers, as detailed later.

### v3

@@ 1. Introduction
Recent advancements in AI (Goodfellow et al., 2014; Li et al., 2022; Touvron et al., 2023; Alayrac et al., 2022) are propelled by a class of models called foundational models (FMs), including large language models (LLMs) and vision-language models (VLMs). FMs leverage billion-scale parameters for improved learning (Yang et al., 2024; Zhang et al., 2024) but impose substantial demands on memory, energy, and compute resources. Recent research has focused on various model compression techniques such as pruning (Frantar and Alistarh, 2023; Ashkboos et al., 2024a) and quantization (Lin et al., 2024a; Shao et al., 2023; Guo et al., 2023; Kang et al., 2024) to reduce memory footprint and computational costs, and make FM inference service more accessible to resource-constrained devices.
Model pruning (Frantar and Alistarh, 2023) reduces memory footprint by removing ineffectual model parameters, such as individual weights (unstructured) or blocks of weights (structured), and storing sparse tensors in a compressed format (Jeong et al., 2023). However, pruning of FMs may be infeasible due to, significant accuracy drops even at low pruning ratios (Kuzmin et al., 2024; Yin et al., [n. d.]) and potential demand for compute and memory-intensive fine-tuning to regain accuracy. Model quantization, on the other hand, reduces the size of a target model by representing weights and/or activations at low precision (Ramachandran et al., 2024b; Kang et al., 2024; Ramachandran et al., 2024a). Recent works on quantization (Dettmers et al., 2022; Guo et al., 2023) have identified that quantizing LLMs is considerably more challenging than quantizing traditional DNNs (Li et al., 2021; Tambe et al., 2020) due to the emergence of large magnitude features known as outliers (Zadeh et al., 2020). These outliers significantly impact model accuracy.
Table 1. Comparison of MicroScopiQ with prior outlier-aware co-design techniques in group A (Zadeh et al., 2020; Park et al., 2018) and B (Guo et al., 2023).
@@ Methods
Techniques in group A, such as OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2023) (algorithm) and GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018), SDQ (Jeong et al., 2024) (architecture co-design) exhibit low accuracy degradation. This is because they typically store outliers at high precision separated from lower precision inliers. However, these techniques result in, (a) low compression factor with high effective bit-width (EBW [^1] The average number of bits used to represent each quantized parameter of a model.) and (b) inefficient hardware and unaligned memory access.
@@
Based on the limitations of existing solutions discussed above, we identify that assignment of higher bit-width for outliers is essential for good accuracy while for aligned memory and hardware efficiency a consistent bit-budget 222number of bits per weight/activation element and data type per tensor element is desired. Here, by consistency we mean that on average each scalar should be represented by a fixed bit-width of a particular data-type. However, these demands are generally conflicting.
@@
Towards achieving this, we present a novel co-design technique for the post-training quantization (PTQ) of FMs, namely MicroScopiQ. Our approach effectively leverages pruning with outlier aware quantization to achieve both memory alignment and improved accuracy. To effectively perform this, for a layer we prune the least important weights based on the Hessian information, and redistribute the additional bits of the outlier weights in the pruned locations. This ensures memory alignment while allowing outlier weights to have higher precision. Additionally, to maintain low error we use the recently proposed MicroScaling (MX) FP data format (Rouhani et al., 2023; Darvish Rouhani et al., 2023) to quantize outlier weights compared to MX-INT inlier quantization. To efficiently support outliers in a different format with different bits at different locations, we present an intelligent NoC architecture called, Redistribution and Coordination NoC (ReCoN). It offers minimal overhead and high throughput outlier processing and reorganization. We then present an accelerator that leverages ReCoN with a simple, homogeneous INT-PE array. Further, we extend the accelerator to be generic enough to support multiple bit-precision (2/4-bit operations). As summarized in Table 1, MicroScopiQ blends the advantages of group A and B techniques while mitigating their specific drawbacks.
@@ 2.1. Model Quantization
$\emph{max}_{\tau}^{b}$ is the maximum representable value of a data format (Ramachandran et al., 2024a). For $b$-bit INT quantization, $\emph{max}_{INT}^{b}=2^{b-1}-1$. After determining the quantization parameters, the quantized tensor is given by (Ramachandran et al., 2024a),
@@
In model quantization, the quantization parameters can be shared at different granularity for different accuracy-overhead trade-offs. In increasing order of overheads, we have per-tensor quantization, wherein the scale factor is shared among all tensor elements. In per-channel quantization, the scale factor is shared per row/column of a tensor. Finally, in group quantization, the parameters are shared at a finer granularity between groups of $k$ (64, 128 etc.) elements in a row or column. These groups are formed by dividing channels into multiple non-overlapping contiguous blocks. In this paper, we adopt MX-INT and MX-FP quantization for inliers and outliers, respectively.
@@ 2.2. Microscaling Data Format
The MX data format proposed by prior works (Rouhani et al., 2023; Darvish Rouhani et al., 2023), is standardized by the Open Compute Project (Project, 2024) with support from Microsoft, Intel, NVIDIA and others. As shown in Figure 1, MX is a variant of block data representation (BDR) (Darvish Rouhani et al., 2023) that defines a format to represent a group of values collectively using shared scale factors. It leverages multi-level, power-of-two scaling, at fine- (level-1, $k_{1}$) and ultra-fine (level-2, $k_{2}$) granularity (Dai et al., 2021; Drumond et al., 2018). The MX data format is characterized by four components: i) scale factors (level-1, 2), ii) data type ($\tau$), iii) bit-width ($b$) and iv) group sizes ($k_{1},k_{2}$). In this paper we denote an MX-FP format as MX-FP-$b$${}_{k_{1},k_{2}}$. In this work, we adopt the version of the MX-FP data format proposed in (Rouhani et al., 2023), employing multi-level scaling. The level-1 scale factor for MX-FP is computed following Equation 1. Conversely, for level-2 scale factor, we identify that MX-FP leverages the sharing of exponent field of FP values (Wang et al., 2018) (referred as $\mu X$ in Figure 1). We show in section 4.2 that by taking advantage of this insight i.e., the concept of shared $\mu X$, we are able to represent FP-outliers in INT format, thereby, enabling the design of simple, homogeneous INT-based PEs. For inliers, we employ MX-INT-$b$${}_{k_{1}}$ with a single level of scale factor following (Project, 2024). This is because, INT format does not possess an exponent field, thereby, a level-2 scale factor similar to MX-FP is not applicable. For simplified understanding, MX-INT-$b$${}_{k_{1}}$ inlier quantization can be viewed as analogous to INT group quantization utilizing an E8M0 scale factor.
@@ 3.1. Limitations of existing techniques
In Table 1, we compare candidate proposals from group A: GOBO (Zadeh et al., 2020) and group B: OliVe (Guo et al., 2023) across various metrics. GOBO is able to achieve high accuracy by retaining outliers at full-precision. It stores outliers separately from low-precision inliers by using sparse representations with the associated outlier indices (see Figure 3(b)). By retaining outliers at full-precision, GOBO results in high EBW. Moreover, the compressed sparse storage and multiple precisions results in unaligned and random memory accesses (Guo et al., 2023), significantly impacting inference latency. Furthermore, GOBO’s outlier handling is hardware inefficient, requiring complex PEs. Similarly, a recent work (Jeong et al., 2024) proposed to decompose a vector of weights into two separate inlier and outlier vectors each quantized in different precisions with outliers at a higher precision.
OliVe (Guo et al., 2023) proposes a scheme to ensure aligned memory access by quantizing inliers and outliers at the same precision (low EBW), but using different data formats. To enable differentiation between the inlier and outlier formats, it prunes the value adjacent to the outlier for use as an identifier (see Figure 3(c)). However, OliVe results in significant accuracy degradation, especially at low precision (see Figure 2(b)), due to: 1) sacrificing a number encoding from inliers for exclusive use as an identifier, reducing the number of representable values in the quantized range, and 2) the rigid assumption of outlier locality–that outliers are never adjacent to each other and only inliers are almost always adjacent to outliers (see section 3.2), leading to unintended outlier pruning. Furthermore, OliVe requires a fairly complex PE design incurring significant encoding/decoding overheads to convert the different formats into a unified processing format (exponent-integer pair). We show that despite quantizing outliers at higher-precision and in a different format, we ensure aligned memory access, simple PE design and minimal hardware overhead.
Figure 2. ((a) Layer-wise distribution of outliers and adjacent outliers as a percentage of total number of weights, (b) Quantization accuracy comparison between OliVe-W4A16 and MicroScopiQ-W2A16 on various benchmarks.
@@ 3.2. Adjacent Outliers Matter
Similar to prior works (Guo et al., 2023; Park et al., 2018), we leverage the $3\sigma$ rule (Pukelsheim, 1994) to categorize weights as outliers. We visually demonstrate the distribution of outliers and adjacent outliers  [^3] We define adjacent outliers as two contiguous outliers along the dot-product dimension (see row 2 of the LLM weight matrix in Figure 3). as a percentage of the total number of weights in a layer across different FMs in Figure 2(a). As the orange box-plot shows, outliers depict a maximum percentage of $\sim 5.1\%$. Outliers are prevalent in FMs, and preserving their values is crucial for maintaining quantized model accuracy. Importantly, from the green box-plots, we observe that modern FMs on average possess $>0.5\%$ adjacent outliers per layer, with some FM layers showing peaks of $>2\%$. This is in stark contrast to the models evaluated by OliVe, such as BERT (Devlin et al., 2018) and OPT (Zhang et al., 2022) which have $<0.04\%$ adjacent outliers (two orders of magnitude lower than FMs like LLaMA3 and LlaVa). This indicates that while pruning values adjacent to outliers could have been ideal for models like BERT (Devlin et al., 2018), it is sub-optimal for most modern FMs as it removes crucial outlier values, leading to higher accuracy degradation. This is evident from Figure 2(b) where OliVe has significant accuracy degradation at 4-bit quantization due to its assumption on outlier locality. Unlike OliVe, MicroScopiQ does not naively prune adjacent values; instead it leverages the Hessian information (Frantar et al., 2022) to identify the least important values to prune, ensuring outlier preservation. This directly translates to high quantized model accuracy and MicroScopiQ at 2-bit consistently outperforms OliVe across different FMs.
@@ 3.3. Outlier Precision and Data Format
The ability of group A techniques like GOBO (Zadeh et al., 2020) to achieve high quantized model accuracy even at extreme quantization levels of inliers ($<4$-bits) is due to retaining outliers at higher precision. This is particularly crucial at ultra-low bit width quantization because, if inliers and outliers are to be quantized to the same precision, there will be higher outlier quantization error due to the reduced representational range. We demonstrate this effect on the MicroScopiQ quantized FM accuracy in Table 6 wherein the quantized FM has poor performance when inliers and outliers are at 2-bits compared to outliers at 4-bits. Furthermore, evidence from recent work (Wu et al., 2023) demonstrates that FP-based formats for LLMs results in superior quantization performance compared to INTs. To validate this, we compare MX-INT v/s MX-FP inlier and outlier quantization in Table 6. Evidently using MX-FP instead of MX-INT for outliers results in better performance. This is due to the higher dynamic range of FPs, which is particularly beneficial at extreme quantization levels. In this work, we quantize outliers at a higher precision (2$\times$) compared to inliers, using MX-FP for outliers and MX-INT for inliers.
@@ 4. MicroScopiQ Quantization Methodology
We present an overview of MicroScopiQ quantization in Figure 3(a) and detail it in Algorithm 1. MicroScopiQ supports various group size granularities and any inlier and outlier (2×2$\times$2 × inliers) data precision. For simplicity, we explain with inlier and outlier precision of 2/4- and 4/8-bit and group sizes of 128 for inliers and 8 for outliers.
Figure 3. (a) Overview of the proposed MicroScopiQ quantization framework depicting methodology of inlier and outlier quantization and redistribution of outlier bits for a sample LLM weight matrix. Comparison against prior quantization frameworks (b) GOBO, and (c) OliVe.
@@ 4.1. Preliminaries
We evaluate the second order derivative of Equation 3, namely the Hessian (Hassibi et al., 1993; LeCun et al., 1989) through Taylor series expansion (Hassibi et al., 1993). Note, the Hessian is the same for all rows depending only on input data and is given as, $\mathbb{H}=2\mathbb{X}\mathbb{X}^{T}$. Its inverse is, $\mathbb{H}^{-1}=(2\mathbb{X}\mathbb{X}^{T}+\lambda\mathbb{I})^{-1}$. We leverage the Hessian information to identify weights with small “saliencies” for pruning (L17 in Algo. 1). The number of inliers to prune is determined by the number of outliers, as detailed later.